Write the virial equation of state .
1
Expert's answer
2019-10-21T07:16:33-0400

Virial equation of state":"

"Z=1+\frac{B}{V}+\frac{C}{V^2}+\frac{D}{V^3}+...."

Where,


"Z=compressibility\ factor\\V=\ molar\ volume\\B,C,D\ and\ so\ on\ are\ virial\ coefficients"